# Science Checker Reloaded: A Bidirectional Paradigm for Transparency and Logical Reasoning

### 21 Feb 2024 v1

@@ III-A Document Retrieval
The system extracts entities and performs expansion from the ontology. We use SciBERT [27] as a model. Each entity constitutes a MUST clause with its expansions as variations. Verbs are augmented with their Wordnet synonyms and will constitute SHOULD clauses. The database indexes the title, abstract, and sections of the article in different fields. The results are ranked by relevance and presented to the user. Depending on the scope of knowledge management, the specificity of the ontology can be of different levels. We used Wikidata [28] as a generalist ontology, and some domain-specific like MeSH [29][30] or internal.
@@ III-B Answer Generation
At various stages of the system, intermediate results are presented to the user to facilitate understanding of the system’s reasoning. These access points are checkpoints for the user, who can thus verify the consistency of the results. Those access points are presented in yellow in the figures 2 and 3. They are the entry points for the user to interact with the system and to check the relevance of the sources and the reasoning of the system.
@@ IV Evaluations
The two blocks of our approach are systems that can be evaluated independently. To evaluate the usability of the whole, it is necessary to carry out user tests [10], especially to evaluate the understanding of the response, the transparency of the system, and the ease of use. However, at the early stage of our work, we conducted quantitative performance evaluations of document retrieval. There is no benchmark sufficiently provided in long document retrieval on scientific data, however the Multilingual Long-Document Retrieval dataset built on Wikipedia and the associated work on BGE M3-Embedding [38] are references to evaluate the performance of our document retrieval system. The dataset consists of text pairs of questions generated by GPT-3.5 based on paragraphs sampled from lengthy articles in 13 languages, and the corresponding articles as the retrieval candidates. We focused on the English language.
We evaluated 3 versions of our system on an Elasticsearch base. The first is the optimized use of the clauses of the base with different types of sub-fields of text, analyzers, and representations in n-grams. The clauses are boosted according to their degree of precision. As there is only one main textual field, the multi-match clauses are evaluated in most_fields. The second is the use of query expansion with an external knowledge graph. We used the Wikidata Ontology [28] to extend the entities detected in the query. These must respond in MUST clauses and boosted. The third is the same version with more flexible clauses on the recognized entities. We report the normalized discounted cumulative gain among the top 10 results.
@@
While the first version gives 1,000 results for each query like the evaluation dataset, it turns out that the two versions with expansion give fewer search results. The absence of results is significant in the version with the expansion of entities executed in MUST clauses, with 18% of queries without relevant documents. This is due to the strong precision constraint on the entities and results in a lower nDCG score than the version without expansion. By relaxing this constraint, with the SHOULD attribution clauses version, we improve the recall on the returned documents. This version, less complex and heavy than the hybrid version of BGE M3, gives performances that are roughly equivalent. In general, our approach, simpler and lighter, gives better performance than dense approaches based on more complex LLMs. As with the second block, its results will need to be refined as the complete system is developed, with additional measures needed to capture the quality of the response, usability, and performance of the complete system in scientific and technical domains.
@@ V Discussion
In a knowledge management context on a specific domain, our retrieval system performs well. However, we observe the limitations of our current system due to the absence of a more general knowledge graph. In these specific cases, the use of the hybrid system proposed by BGE M3 gives better results, despite these results being frozen by the model’s weights. The development of top-level ontologies that allow the interoperability of more specific graphs, such as the WikiProject Ontology, is a method to have a system that improves with the advancement of knowledge. This is a development path for our system to have better knowledge coverage, in addition to regularly updated article databases.

### 14 Mar 2024 v2

@@ III-A Document Retrieval
The system extracts entities and performs expansion from the ontology. We use SciBERT [27] as a model. Each entity constitutes a MUST clause with its expansions as variations. Verbs are augmented with their Wordnet synonyms and will constitute SHOULD clauses. The database indexes the title, abstract, and sections of the article in different fields. The results are ranked by relevance and presented to the user. Depending on the scope of knowledge management, the specificity of the ontology can be of different levels. We used Wikidata [28] as a generalist ontology, and some domain-specific ontologies like MeSH [29][30] or internal ontology.
@@ III-B Answer Generation
At various stages of the system, intermediate results are presented to the user to facilitate understanding of the system’s reasoning. These access points are checkpoints for the user, who can thus verify the consistency of the results. Those access points are presented in yellow in Figures 2 and 3. They are the entry points for the user to interact with the system and to check the relevance of the sources and the reasoning of the system.
@@ IV Evaluations
The two blocks of our approach are systems that can be evaluated independently. To evaluate the usability of the whole, it is necessary to carry out user tests [10], especially to evaluate the understanding of the response, the transparency of the system, and the ease of use. However, at the early stage of our work, we conducted quantitative performance evaluations of document retrieval. There is no benchmark sufficiently provided in long document retrieval on scientific data, however the Multilingual Long-Document Retrieval dataset built on Wikipedia and the associated work on M3-Embedding [38] (Multi-Lingual, Multi-Functionality, Multi-Granularity Text Embeddings) are references to evaluate the performance of our document retrieval system. The dataset consists of text pairs of questions generated by GPT-3.5 (Generative Pre-trained Transformer) based on paragraphs sampled from lengthy articles in 13 languages, and the corresponding articles as the retrieval candidates. We focused on the English language.
We evaluated 3 versions of our system on an Elasticsearch base. The first is the optimized use of the clauses of the base with different types of sub-fields of text, analyzers, and representations in n-grams. The clauses are boosted according to their degree of precision. As there is only one main textual field, the multi-match clauses are evaluated in most_fields. The second is the use of query expansion with an external knowledge graph. We used the Wikidata Ontology [28] to extend the entities detected in the query. These must respond in MUST clauses and boosted. The third is the same version with more flexible clauses on the recognized entities. We report the normalized Discounted Cumulative Gain (nDCG) among the top 10 results.
@@
While the first version gives 1,000 results for each query like the evaluation dataset, it turns out that the two versions with expansion give fewer search results. The absence of results is significant in the version with the expansion of entities executed in MUST clauses, with 18% of queries without relevant documents. This is due to the strong precision constraint on the entities and results in a lower nDCG score than the version without expansion. By relaxing this constraint, with the SHOULD attribution clauses version, we improve the recall on the returned documents. This version, less complex and heavy than the hybrid version of M3-Embedding, gives performances that are roughly equivalent. In general, our approach, simpler and lighter, gives better performance than dense approaches based on more complex LLMs. As with the second block, its results will need to be refined as the complete system is developed, with additional measures needed to capture the quality of the response, usability, and performance of the complete system in scientific and technical domains.
@@ V Discussion
In a knowledge management context on a specific domain, our retrieval system performs well. However, we observe the limitations of our current system due to the absence of a more general knowledge graph. In these specific cases, the use of the hybrid system proposed by M3-Embedding gives better results, despite these results being frozen by the model’s weights. The development of top-level ontologies that allow the interoperability of more specific graphs, such as the WikiProject Ontology, is a method to have a system that improves with the advancement of knowledge. This is a development path for our system to have better knowledge coverage, in addition to regularly updated article databases.